Dylan computed a statistical power for his experiment to be 0.7. What is the value of Beta (B)? 0.3? 0.7? 70%? 0.49?
Answer #1

The statistical power for Dylan’s experiment is 0.7.

The value of beta \left (\beta \right ) can be obtained as,

\begin{array}{c} Power = 1 - \beta \\ 0.7 = 1 - \beta \\ \beta = 1 - 0.7\\ \beta = 0.3 \end{array}

Therefore, the value of beta is 0.3.